# Theories of “Sexuality” in Natural Language Processing Bias Research

### v1

@@ Abstract
Jacob T. Hobbs Department of Computer Science Department of Women, Gender, and Sexuality University of Virginia jacobhobbs@virginia.edu
@@ 1 Introduction
In an effort to mitigate these harms, sexuality is increasingly becoming a common demographic axis for which to measure bias in NLP systems. Building on seminal work uncovering algorithmic racial and gender bias (Bolukbasi et al., 2016; Caliskan et al., 2017; Manzini et al., 2019; Sun et al., 2019), a handful of authors have begun to research how queer sexualities are processed in various NLP tasks such as language generation (Sheng et al., 2019; Dhingra et al., 2023; Felkner et al., 2024), machine translation (Stewart and Mihalcea, 2024), sentiment analysis (Ungless et al., 2023), hate speech/toxicity detection (Davani et al., 2023), and even heteronormativity detection (Vásquez et al., 2022). Scholars in the field of gender and sexuality studies have noted, however, that a stable definition of queerness is slippery at best, owing to the social construction of sexuality and its entanglement with constricting systems of power (Foucault, 1978; Butler, 1990). Therefore, we are primarily interested in how technical researchers articulate and codify sexuality, centering our analysis on NLP bias research.
@@ 2.2 Sexuality, Gender, and Social Theory
A single, stable definition of sexuality is inherently illusive. From biology to sociology, theology to sexology, feminist studies to queer theory, there are vast differences between how sexuality is researched and theorized across disciplines. As such, for the purposes of this study it would be improper to narrowly confine "sexuality" under a single all-encompassing definition. Therefore, we briefly outline how sexuality can be conceptualized along various continua, how gender and race are inextricably and historically linked to modern ideas of sexuality, and how queer and feminist theories may further trouble a stable definition of sexuality.
@@
A useful heuristic for understanding sexuality is to place an individual’s sexual and romantic preferences (or non-preferences) on various continua. One of the earliest attempts to categorize sexuality in this way was "The Kinsey Scale," which saw sexuality as a variable mixture of homosexual and heterosexual urges (Kinsey et al., 1948). Moving slightly past finite identity categories, The Kinsey Scale allowed medical experts to categorize sexuality beyond a binary (homosexuality/heterosexuality) or trinary (homo/hetero/bisexuality) model. Similar methods of placing sexuality on a scale or spectrum are still popular today. Someone who feels a strong alignment with a certain sexual or romantic identity may choose to self-identify their sexual/romantic orientation, with common identities falling under broader homosexual/romantic, heterosexual/romantic, or bisexual/romantic categories. Asexuality or aromanticism are identities that describe less or no sexual/romantic attraction, and many see the "aro/ace" spectrum as another component of any person’s sexuality. A somewhat comprehensive list of commonly used identities can be found on websites like the community-curated LGBTQIA+ Wiki [^1] https://lgbtqia.fandom.com/wiki/.
@@ 3.1 Method
Sexuality Focus Is measuring sexuality bias the primary focus of the article
@@
Intersectionality Is sexuality bias measured together with other oppressions
Language What language(s) are investigated
Technology What technology is examined
@@ 3.2 Results
Finally, Table 5, 6, and 7 report the Sexuality Focus, Beyond Duality, and Intersectionality classifications, respectively. We find that a majority of papers do not have a sexuality focus or utilize an intersectional framework, meaning most papers consider multiple demographics but do so independently in their bias analysis. Additionally, we find that roughly half of the papers (25, or 45.5%), go beyond a simple heterosexual/LGBTQ+ binary comparison structure. This means an article may consider multiple queer identities (and breaks down results by identity) or simply does not consider heterosexuality (common when only considering "marginalized" or "protected" groups).
@@
many identities considers >3 sexuality identities 27
@@
spectrum acknowledges sexuality is a spectrum/complex 6
undefined no clear framework 3
underspecified does not explicitly/clearly define sexuality 38
@@
identity word list uses a set list of identities 40
@@
multiple explicitly considers many dimensions of bias 2
occupation uses occupation titles as a measure of bias 5
QA asks: does a QA model prefer one answer choice over another? 6
regard score comparisons: uses "regard" (Sheng et al., 2019) 5
translation measures machine translation accuracy 2
word embeddings compares sexuality identity word vectors 7
@@
Sexuality Focus #
no 47
yes 8
Table 5: Is measuring sexuality-based bias in NLP systems the primary focus of the article?
Beyond Duality #
no 30
yes 25
Table 6: Does the article go beyond a simple heterosexual/LGBTQ+ binary comparison structure?
Intersectional #
no 48
yes 7
Table 7: Is sexuality bias measured together with other oppressions? The paper must consider multiple oppressed identities simultaneously. It is not sufficient to analyze multiple biases separately.
@@ 4.1 Gender/Sexuality Conflation
The most basic type of gender/sexuality conflation is when an article classifies sexual identities as gender identities or gender identities as sexual identities. Mukherjee et al. (2023), for example, does the latter when they characterize the type of bias "LGBTQ+" people face simply as "sexuality bias," grouping transgender people into specifically sexuality-based bias evaluations. Hassan et al. (2021), on the other hand, categorizes the words "…lesbian, gay, bisexual, [and] asexual…" under the category of "Gender Identity," demonstrating a gross misunderstanding of both sexuality and gender.
@@ 4.3 Heteronormativity
Nangia et al. (2020) frames white as an appropriate "contrasting advantaged group" to lesbians, a "historically disadvantaged group." Beyond the issue that white is a racial identity that should not be used to test for sexuality bias, the assertion that white, rather than straight, is the most logical contrasting word to lesbian demonstrates that heterosexuality is thoroughly ingrained as the default in US culture, so much so that the annotator for this particular example may have perceived whiteness and heterosexuality as overlapping or mutually reinforcing categories of privilege.
Counterfactuals have the potential to go beyond centering heterosexuality, but no paper that uses this structure attempts to do so. Felkner et al. (2023) employ contrasting pairs and include myriad queer identities, but, similar to Nangia et al. (2020), always use heterosexuality as the default with which to compare. We believe researchers can uncover more nuanced intragroup biases by de-centering heterosexuality and directly comparing queer sexualities to one another, which is discussed further in Section 5.4.
@@ 5.4 Decenter Heterosexuality as the Default
As discussed in Section 4.3, many papers surveyed use methodologies that center heterosexuality as the default. In fact, heterosexuality is often not operationalized as a sexual identity, evidenced by the handful of papers surveyed that claim to analyze sexuality-based bias but only consider queer sexualities. While having a clear theory of sexuality, as discussed above, may help address this issue, we observe that heteronormativity can still permeate a paper’s methodology. For instance, WinoQueer (Felkner et al., 2023) encouragingly makes clear their theorization of sexuality and reports detailed results broken down by identity category. However, their methodology is based off CrowS-Pairs (Nangia et al., 2020), which employ counterfactual stereotypical/anti-stereotypical pairs to evaluate bias. While WinoQueer is able to show clear model bias by comparing myriad queer identities to "a corresponding non-LGBTQ+ identity," this approach is constrained by its assumption that queer identities in fact have "corresponding" opposites.
We ask, then, what nuanced biases could be uncovered if we don’t take heterosexuality as the baseline for which to compare? If a counterfactual data structure must be used, could two queer identities be compared, not as stereotypical opposites but as a more privileged / more marginalized pair? In other words, if we decenter heterosexuality, what subtle intragroup biases could we uncover? Though contrasting pairs that reveal queer/non-queer model preferences are useful, how could future work that completely removes heterosexuality from the equation center queerness in bias evaluations? We discuss potential strategies for this in the following section.
@@ 6 Conclusion
Through a survey of 55 papers that concern sexuality bias in Natural Language Processing, we found that a majority of NLP researchers do not clearly articulate what sexuality is or why particular proxies for sexuality should be used. Further, we found that few papers operationalize nuanced or intersectional methodologies that account for the specific realities of those who may face the greatest systematic bias from NLP systems. Finally, we proposed actionable solutions that encourage the NLP community to make their conception of sexuality explicit and to consider more complex frameworks grounded in queer experiences and informed by interdisciplinary scholarship.
@@ Appendix B Additional Survey Results
Additional information collected about all surveyed papers, including what language the paper investigated and what technology was examined, is available in Table 8 and 9.

### v2

@@ Abstract
Jacob T. Hobbs Department of Computer Science Department of Women, Gender, and Sexuality University of Virginia fdu5ff@virginia.edu
@@ 1 Introduction
In an effort to mitigate these harms, sexuality is increasingly becoming a common demographic axis for which to measure bias in NLP systems. Building on seminal work uncovering algorithmic racial and gender bias (Bolukbasi et al., 2016; Caliskan et al., 2017; Manzini et al., 2019; Sun et al., 2019), a handful of authors have begun to research how queer sexualities are processed in various NLP tasks such as language generation (Sheng et al., 2019; Dhingra et al., 2023; Felkner et al., 2024), machine translation (Stewart and Mihalcea, 2024), sentiment analysis (Ungless et al., 2023), hate speech/toxicity detection (Davani et al., 2023), and even heteronormativity detection (Vásquez et al., 2022). Scholars in the field of gender and sexuality studies have noted, however, that a stable definition of queerness is slippery at best, owing to the social construction of sexuality and its entanglement with constricting systems of power (Foucault, 1978; Butler, 1990). Butler in particular emphasizes how language itself plays an important role in constructing identity categories in the first place. Therefore, we are primarily interested in how technical researchers articulate, codify, and thus reinforce various theories of sexuality, centering our analysis on NLP bias research.
@@ 2.2 Sexuality, Gender, and Social Theory
A single, stable definition of sexuality is inherently illusive. From biology to sociology, theology to sexology, feminist studies to queer theory, there are vast differences between how sexuality is researched and theorized across disciplines. The very notion of sexuality as an identity characteristic, rather than a behavioral tendency, didn’t take root in popular culture until the early 20th century (Katz, 2007). As such, for the purposes of this study it would be improper to narrowly confine "sexuality" under a single all-encompassing definition. Therefore, we briefly outline how sexuality can be conceptualized along various continua, how gender and race are inextricably and historically linked to modern ideas of sexuality, and how queer and feminist theories may further trouble a stable definition of sexuality.
@@
A useful heuristic for understanding modern sexuality is to place an individual’s sexual and romantic preferences (or non-preferences) on various continua. One of the earliest attempts to categorize sexuality in this way was "The Kinsey Scale," which saw sexuality as a variable mixture of homosexual and heterosexual urges (Kinsey et al., 1948). Moving slightly past finite identity categories, The Kinsey Scale allowed medical experts to categorize sexuality beyond a binary (homosexuality/heterosexuality) or trinary (homo/hetero/bisexuality) model. Similar methods of placing sexuality on a scale or spectrum are still popular today. Someone who feels a strong alignment with a certain sexual or romantic identity may choose to self-identify their sexual/romantic orientation, with common identities falling under broader homosexual/romantic, heterosexual/romantic, or bisexual/romantic categories. Asexuality or aromanticism are identities that describe less or no sexual/romantic attraction, and many see the "aro/ace" spectrum as another component of any person’s sexuality. A somewhat comprehensive list of commonly used identities can be found on websites like the community-curated LGBTQIA+ Wiki [^1] https://lgbtqia.fandom.com/wiki/.
@@ 3.1 Method
Sexuality Focus Is measuring sexuality bias the primary focus of the article?
@@
Intersectionality Is sexuality bias measured together with other oppressions?
Language What language(s) are investigated?
Technology What technology is examined?
@@ 3.2 Results
Finally, Figure 1 reports the Sexuality Focus, Beyond Duality, and Intersectionality classifications. We find that a majority of papers do not have a sexuality focus or utilize an intersectional framework, meaning most papers consider multiple demographics but do so independently in their bias analysis. Additionally, we find that roughly half of the papers (25, or 45.5%), go beyond a simple heterosexual/LGBTQ+ binary comparison structure. This means an article may consider multiple queer identities (and breaks down results by identity) or simply does not consider heterosexuality (common when only considering "marginalized" or "protected" groups).
@@
underspecified does not explicitly/clearly define sexuality 38
many identities considers >3 sexuality identities 27
@@
undefined no clear framework 3
@@
identity word list uses a set list of identities 40
@@
word embeddings compares sexuality identity word vectors 7
QA asks: does a QA model prefer one answer choice over another? 6
occupation uses occupation titles as a measure of bias 5
regard score comparisons: uses "regard" (Sheng et al., 2019) 5
multiple explicitly considers many dimensions of bias 2
translation measures machine translation accuracy 2
@@
Figure 1: Additional paper classifications. Intersectional: Is sexuality bias measured together with other oppressions? The paper must consider multiple oppressed identities simultaneously. Sexuality Focus: Is measuring sexuality-based bias in NLP systems the primary focus of the article? Beyond Duality: Does the article go beyond a simple heterosexual/LGBTQ+ binary comparison structure?
@@ 4.1 Gender/Sexuality Conflation
The most basic type of gender/sexuality conflation is when an article classifies sexual identities as gender identities or gender identities as sexual identities. Mukherjee et al. (2023), for example, does the latter when they characterize the type of bias "LGBTQ+" people face simply as "sexuality bias," grouping transgender people into specifically sexuality-based bias evaluations. Hassan et al. (2021), on the other hand, categorizes the words "…lesbian, gay, bisexual, [and] asexual…" under the category of "Gender Identity."
@@ 4.3 Heteronormativity
Nangia et al. (2020) frames white as an appropriate "contrasting advantaged group" to lesbians, a "historically disadvantaged group." Beyond the issue that white is a racial identity that should not be used to test for sexuality bias, the assertion that white, rather than straight, is the most logical contrasting word to lesbian demonstrates that heterosexuality is thoroughly ingrained as the default in US culture, so much so that whiteness and heterosexuality are implicitly theorized as overlapping or mutually reinforcing categories of privilege.
Counterfactuals may have the potential to go beyond centering heterosexuality, but no paper that uses this structure attempts to do so. Felkner et al. (2023) employ contrasting pairs and include myriad queer identities, but, similar to Nangia et al. (2020), always use heterosexuality as the default with which to compare. We believe researchers can uncover more nuanced model bias by decentering heterosexuality and moving beyond direct comparison of pre-defined identity keywords, which is discussed more in Secion 5.5.
@@ 5.4 Decenter Heterosexuality as the Default
As discussed in Section 4.3, many papers surveyed use methodologies that center heterosexuality as the default. In fact, heterosexuality is often not operationalized as a sexual identity, evidenced by the handful of papers surveyed that claim to analyze sexuality-based bias but only consider queer sexualities. While having a clear theory of sexuality, as discussed above, may help address this issue, we observe that heteronormativity can still permeate a paper’s methodology. For instance, WinoQueer (Felkner et al., 2023) encouragingly makes clear their theorization of sexuality and reports detailed results broken down by identity category. However, their methodology is based off CrowS-Pairs (Nangia et al., 2020), which employ counterfactual stereotypical/anti-stereotypical pairs to evaluate bias. While WinoQueer is able to show clear model bias by comparing myriad queer identities to "a corresponding non-LGBTQ+ identity," this approach is constrained by its assumption that queer identities in fact have "corresponding," mutually exclusive opposites. We ask, then, what nuanced biases could be uncovered if we don’t take heterosexuality as the baseline for which to compare? We discuss potential strategies for this in the following section.
@@ 6 Conclusion
Through a survey of 55 papers that concern sexuality bias in Natural Language Processing, we found that a majority of NLP researchers do not clearly articulate what sexuality is or why particular proxies for sexuality should be used. Further, we found that few papers operationalize nuanced or intersectional methodologies that account for the specific realities of those who may face the greatest systematic bias from NLP systems. We proposed multiple actionable recommendations, distilled into the following checklist for NLP bias researchers:
  • Make theory of sexuality explicit, including why certain identities are included/excluded.
  • Refrain from treating queerness as a monolith; adopt intersectional methodologies and make precise statements about exactly who is harmed by a biased system.
  • Engage with relevant interdisciplinary literature, community members, and stakeholders throughout the research process.
Finally, we encourage sociotechnical research that aims to "queer" NLP methodologies by challenging fundamental assumptions about identity and bias.
@@ Appendix B Additional Survey Results
Additional information collected about all surveyed papers, including what language the paper investigated and what technology was examined, is available in Table 5 and 6.